OpenStudy (anonymous):

(800 radical2)((radical3)/2)

OpenStudy (anonymous):

\[\frac{ 800\sqrt{2}*\sqrt{3} }{ 2 }=?\] when the numbers under the radical are not negative they can be multiplied.

OpenStudy (anonymous):

if both are negative they can't be multiplied.